OpenStudy (anonymous):

Can someone help me on this algorithm question? I don't understand what the problem is asking for... Suppose a parallel computer system has one thousand processors which are programmed to solve a particular problem in parallel. Derive a reasonable estimate for the largest values of n this computer system can solve in 1 day, 1 month, 1 year, and 10 years, respectively, assuming the algorithms for solving this problem of size n requires O(n), O(n^2), O(n^3) and O(2^n) instructions such that one instruction takes 1 nanoseconds.

ganeshie8 (ganeshie8):

question is asking u to calculate how many instructions can be executed

ganeshie8 (ganeshie8):

\(1 \ day = 24 \times 60 \times 60 \times 10^9 \ ns\)

ganeshie8 (ganeshie8):

I'll solve them for 1 day, you may solve them for the rest for a problem of size O(n) to finish in 1day, the size cannot be larger than \( 24 \times 60 \times 60 \times 10^9\) similarly, for a problem of size O(n^2) to finish in 1day, the size cannot be larger than \( (24 \times 60 \times 60 \times 10^9)^{1/2}\) for a problem of size O(n^3) to finish in 1day, the size cannot be larger than \( (24 \times 60 \times 60 \times 10^9)^{1/3}\) for a problem of size O(2^n) to finish in 1day, the size cannot be larger than \( \log_2 (24 \times 60 \times 60 \times 10^9)\)

ganeshie8 (ganeshie8):

see if those make some sense :)

OpenStudy (anonymous):

yea, thx. i just got it after reading it 30 times over haha.

ganeshie8 (ganeshie8):

good :)